# CoBOS: Constraint-Based Online Scheduler for Human-Robot Collaboration

### v1

@@ VIII Conclusion
In our comprehensive evaluation, we compare our proposed online constraint-based scheduling method against three baseline methods. These comparisons span across seven classes of tasks designed to test different aspects of uncertainty and decision-making flexibility in human-robot collaboration in assembly processes.
@@
Our method significantly outperforms all baseline methods, particularly in the more complex task classes. This demonstrates its ability to manage uncertainties and enhance human-robot collaboration in assembly tasks. We gave a preview of our proof-of-concept implementation on a real Franka Emika Panda robot.
@@
The developed tools and a clear baseline enable further research regarding uncertainty-aware scheduling for collaborative manufacturing. Specifically, how to enable the CoBOS to take known distributions into account, systematic experiments with the real setup, the introduction of more complex tasks, and the learning of probabilistic distributions from experience.

### v2

@@ VIII Conclusion
In our comprehensive evaluation, we compare our proposed online constraint-based scheduling method against two baseline methods and a method from the literature. These comparisons span across seven classes of tasks designed to test different aspects of uncertainty and decision-making flexibility in human-robot collaboration in assembly processes.
@@
Our method significantly outperforms the other methods, particularly in the more complex task classes. This demonstrates its ability to manage uncertainties and enhance human-robot collaboration in assembly tasks. We gave a preview of our proof-of-concept implementation on a real Franka Emika Panda robot.
@@
The developed tools and a clear baseline enable further research regarding uncertainty-aware scheduling for collaborative manufacturing. Specifically, how to enable CoBOS to take known distributions into account, systematic experiments with the real setup, the introduction of more complex tasks, and the learning of probabilistic distributions from experience.